# Signed Networks: theory, methods, and applications

### v1

@@ 4 Network structure and inference
Historically, the approach to the study of networks has been two-fold: on one hand, we are interested in investigating their complex structures, while on the other hand, we might want to study their evolution over time and the dynamical processes occurring on them [BOCCALETTI2006]. These two aspects are strongly interconnected, as the structure of a network can influence its dynamics (e.g., opinions will spread differently in a highly clustered network than in a non-clustered one) and vice versa (e.g., a change in opinion might break a link or alter its sign).
@@ 4.1 Null models
In signed networks, the most common use of null models is to detect the presence of strong or weak structural balance in empirical networks. Real-world networks often deviate from a state of perfect balance; therefore, a statistical analysis is required to determine the significance of specific patterns compared to a null model. Part of the debate over whether social systems exhibit strong or weak structural balance comes from conflicting results caused by differences in the null models employed [maozWhatEnemyMy2007, kirkleyBalanceSignedNetworks2019]. In what follows, we present null models that have mostly been developed for assessing balance. However, null models also play an important role in other contexts, for instance, to predict the creation of enmities, detect communities, or assess the importance of homophily and heterophily in signed networks, and can be used in such cases, beyond structural balance. For additional details on how to choose appropriate null models, see [hobsonGuideChoosingImplementing2021].
Figure 9: Null models for signed networks. The figure shows the different null models that can be used in signed networks, depending on the constraints that are preserved. The top row shows simple re-wiring or sign shuffle null models, while the bottom row depicts two models that preserve the signed degree. The asterisk (*) indicates that the signed degree in the STP model is preserved on average. Models in the right column preserve the underlying topology. All models preserve the ratio of negative edges in the network. This figure is an adaptation of Figure 2 from [haoProperNetworkRandomization2023].
Following the classification presented in [haoProperNetworkRandomization2023], null models can be organised by the constraints they preserve, such as topology, node degree, or signed degree. Figure 9 illustrates examples of the different null models presented below. In addition to the classification by Hao et al. [haoProperNetworkRandomization2023], we also consider null models based on conditional probability methods, stochastic block models, and maximum-entropy formulations.
@@
This model disrupts topology while maintaining the underlying degree (i.e. the number of neighbours of each node, regardless of sign). One way to implement this null model is to randomly select two edges and interchange their source or target nodes with equal probability. To avoid multi-edges, this change is only accepted if the newly defined edge does not exist in the network. While this model is widely used in unsigned networks because it preserves node degree, it is less suitable for signed networks, because it disrupts the heterogeneity of nodes’ signed degrees and may conceal patterns related to the distribution of positive and negative links.
Sign shuffle model [szellMultirelationalOrganizationLargescale2010, leskovecSignedNetworksSocial2010]
@@
Here, the topology remains fixed while the signs of the edges are randomly redistributed. An analogous approach, shown to provide similar results in terms of balance [facchettiComputingGlobalStructural2011], is to draw the edge signs as independent and identically distributed variables from a Bernoulli distribution. While this model preserves the total ratio of positive versus negative edges, it assumes a homogeneity in the individual ratios of each node (i.e., it assumes all nodes have the same proportion of positive edges). This assumption is dangerous because in most real-life networks there is a clear heterogeneity between individuals, some of them being more “friendly” than others [haoProperNetworkRandomization2023] (see Section 5.2.3). Moreover, in [facchettiComputingGlobalStructural2011] it is discussed how signed degree homogeneity generates more unbalanced networks and thus results in conclusions in favour of balance theory when comparing real data with such models. The explanation for this is that nodes with a lot of enmities are easier to “cast out” than uniformly distributed conflicts in the system. Another issue with this null model, brought forward by [fengTestingBalanceSocial2022], is that it assumes that positive and negative edges are interchangeable, therefore testing a stronger statement than just balance. To solve this issue, the authors propose a modified permutation-based null model that stratifies the uniform permutation to edges with the same level of embeddedness.
Signed re-wire model [liConstructingRefinedNull2021]
@@
This model follows the same logic as its unsigned counterpart, explained above. However, to maintain the signed degree, it rewires the positive and negative edges separately. An equivalent implementation is to “cut” all edges in half and reattach them randomly, as long as the two halves that are reattached have the same sign [^17] In the case of a directed network, the reattachment must also follow the direction of the edge.. This approach was used, for example, in the evaluation of balance in ecological networks [saizEvidenceStructuralBalance2017]. An alternative model, proposed in [derrSignedNetworkModeling2018], is the signed Chung-Lu model, which preserves the degrees of the nodes on average rather than exactly. The authors employ this model to generate synthetic signed networks that reproduce both the global sign distribution and the proportion of balanced triangles observed in real data.
Signed Degree and Topology Preserving (STP) Model [haoProperNetworkRandomization2023]
@@
Hao et al. [haoProperNetworkRandomization2023] introduced a model that preserves both the network topology and the signed node degree. It is based on the maximum entropy framework [squartiniAnalyticalMaximumlikelihoodMethod2011] and provides the exact probabilities of occurrence of random graphs with both constraints. Note that, while it does preserve the signed node degree on average, it does not preserve the exact, individual signed degrees for each node.
Conditional-probability approach [lernerStructuralBalanceSigned2016]
@@
Lerner et al. [lernerStructuralBalanceSigned2016] proposed a different approach in terms of conditional probabilities, specifically the probability of a link having a particular sign given the presence of the link itself, instead of marginal probabilities. This approach, while it has a different implementation, has a similar conceptualisation to the topology-preserving models. The decomposition of the marginal probability of a negative link, then, would be given by:
@@
Stochastic block models [jiangStochasticBlockModel2015, Yang_Zhao_Liu_2015, galloAssessingFrustrationRealworld2024]
A common null model for community detection is the stochastic block model (SBM). It assumes that nodes can be grouped into latent blocks and that the probability of observing a positive or negative edge depends only on the blocks of the endpoints. The model parameters $p^{+}_{rs}$ and $p^{-}_{rs}$ form a block matrix that specifies the probabilities of positive and negative links between nodes belonging to blocks $r$ and $s$, respectively. Jiang et al. [jiangStochasticBlockModel2015] proposed a generalised signed SBM where they differentiate positive and negative block matrices. Yang et al. [Yang_Zhao_Liu_2015, yangCommunityMiningSigned2007] then developed a Bayesian signed SBM which explicitly accounted for edge density and frustration in the generative model. Finally, Gallo et al. [galloAssessingFrustrationRealworld2024] derived the signed SBM from an exponential random graph framework and connected it to the theory of relaxed balance (see Section 3.4).
Maximum-entropy approaches [galloTestingStructuralBalance2024, becattiEntropybasedRandomizationRating2019, hao2025social, galloPatternsLinkReciprocity2024]
Gallo et al. [galloTestingStructuralBalance2024] have extended exponential random graph models to signed networks with both global and local constraints. Their global Signed Configuration Model with Fixed Topology (SCM-FT) imposes the topology as a hard constraint, while the positive and negative degrees of each node are matched only on average, as in the STP model explained above. The authors also define three additional variants, aligned with the types of constraints explained above: a homogeneous model with varying topology (Signed Random Graph Model or SRGM), with the same constraints as the Re-wire model, a homogeneous model with fixed topology (SRGM -FT), similar to the Sign shuffle model, and a heterogeneous model with varying topology (SCM), similar to the Signed re-wire model. A bipartite version of the SCM can also be recovered from the Bipartite Score Configuration Model in [becattiEntropybasedRandomizationRating2019], and extensions for directed signed networks can be found in [hao2025social] and [galloPatternsLinkReciprocity2024].
@@
The choice of a null model strongly shapes the interpretation of structural balance and other patterns in signed networks. In [haoProperNetworkRandomization2023], the authors argue that matching the signed degrees is a relevant and often forgotten feature, while [lernerStructuralBalanceSigned2016] shows that results may support or contradict a hypothesis of balance depending on whether conditional or marginal probabilities of edge signs are considered. Similarly, [galloTestingStructuralBalance2024] finds that conclusions about strong or weak balance depend on whether heterogeneous sign propensities are included and on whether the network topology is preserved.
Beyond testing balance, null models serve broader purposes such as identifying mesoscale organisation, evaluating link or sign prediction, or constructing bipartite projections [gallo2025statistically]. Their design should always reflect the constraints of the system under study. Re-wire models are useful in scenarios where individuals are free to decide who and how to interact with other individuals (e.g., an unconstrained social system like social media). If users’ friendliness and hostility are homogeneous, it is safe to use models that do not preserve signed degrees. However, if there are different tendencies in the way users interact with each other, the signed re-wire model would be more appropriate. Conversely, if the individuals cannot choose who they interact with but they choose the how (e.g., a high school), we would prefer topology-preserving models like a sign shuffle model, or models such as STP or SCM-FT. Ideally, one can go a few steps further and design a tailored null model when information on the allowed connections is available, since generally the topology of a network is not completely arbitrary nor strongly determined. In that case, these constraints could be incorporated in maximum-entropy-based models similar to the STP or SCM-FT [galloTestingStructuralBalance2024].
@@ 4.2 Node embeddings
Node embedding methods are a powerful branch of the representation learning field, used to learn compact, low-dimensional representations of nodes, such that proximity in the embedding space reflects the underlying network structure. In this section, we review the methods specifically developed for signed networks, with a summary provided in Table LABEL:tab:node-embeddings. For a more complete overview of the field, we refer to a recent survey [zhang2024signed].
The core principle underlying many of these methods is synthesised in the title of one of the first papers in this domain: When Everybody Can Sit Closer to Friends than Enemies [kermarrecSignedGraphEmbedding2014]. The goal of these methods is to embed the nodes in a space, such that the distance between positively connected nodes (friends) is smaller than between negatively connected nodes (enemies). However, in general graphs, this ideal positioning cannot be achieved, as Kermarrec and Thraves [kermarrecSignedGraphEmbedding2014] proved even for simple graphs. Consequently, most algorithms relax this objective and instead aim to minimise the number of violated constraints by optimising equivalent spectral, probabilistic, or geometric criteria. Early work in this domain focused on Euclidean and spectral approaches, while more recent contributions have leveraged advances in natural language processing, probabilistic modelling, hyperbolic geometry, and graph neural networks (GNNs).
One of the first contributions to node embeddings for signed networks was proposed by Kunegis et al. [kunegisSpectralAnalysisSigned2010], who introduced a spectral approach to represent nodes in a $k$-dimensional Euclidean space. In a nutshell, the eigenvectors corresponding to the $k$ smallest eigenvalues of the opposing Laplacian are used to embed nodes in $\mathbb{R}^{k}$, so that Euclidean proximity reflects many positive links and few negative ones. This spectral embedding is particularly useful for community detection in signed graphs, as we discuss in Section 4.4. A related Euclidean embedding proposed in [diazdiazGremban] leverages the fact that the communicability distance is Euclidean to embed the nodes within a hypersphere.
Another spectral embedding technique, inspired by physical arguments, is the SHEEP algorithm by Babul and Lambiotte [babulSHEEPSignedHamiltonian2024]. The method is based on a Hamiltonian where positive edges act as springs pulling nodes together and negative edges act as repulsive forces.
@@
Minimising this Hamiltonian under a spherical constraint reduces to an eigenvalue problem involving the repelling Laplacian (Eq. 2.3). The method has been successfully applied to empirical data, including detecting opinion polarisation in online social networks [candellone2025negative].
A different family of methods takes inspiration from word embedding techniques like Mikolov’s word2vec model [mikolovEfficientEstimationWord2013] and adapts them to networks. In word2vec, words are mapped to vectors by scanning sentences with sliding windows and adjusting the vectors so that words appearing in similar contexts lie close together. Grover and Leskovec adapted this idea to graphs with node2vec [groverNode2vecScalableFeature2016], replacing sentences by random walks on the network and words by nodes. Nodes that frequently co-occur along walks thus become close in the embedding space. The first methods to extend this principle to signed graphs were SNE [yuanSNESignedNetwork2017] and Sign2vec  [islamDistributedRepresentationsSigned2018]. In these works, walks are adapted to incorporate edge signs, so that the learned vectors reflect both proximity and polarity. More recently, the stance embeddings model (SEM[pougue-biyongLearningStanceEmbeddings2022] added edge attributes, enabling embeddings that account for topical differences. Alternatively, one can map the signed network into an unsigned one without losing the information carried by negative edges. This approach, used in pipelines like ROSE [javariROSERolebasedSigned2020], allows the direct use of methods such as node2vec while retaining the information contained in the negative edges in the embedding. Other approaches to signed network embedding are based on correlations between signed random walks [huangPOLEPolarizedEmbedding2022], Bayesian approaches [maraCSNEConditionalSigned2020], hyperbolic geometry [songHyperbolicNodeEmbedding2021], or latent distance models that structure embeddings within polytopes [nakisCharacterizingPolarizationSocial2023, nakisHybridMembershipLatent2023].
The introduction of Graph Neural Networks (GNNs) opened a new phase for representation learning on signed graphs and extended node-embedding ideas into deep models. One of the earliest attempts in this direction is SiNE [wangSignedNetworkEmbedding2017], a feedforward neural network that learns embeddings by enforcing friends to be closer than foes. This line of work has evolved toward graph convolutional neural networks [derrSignedGraphConvolutional2018, kimTrustworthinessDrivenGraphConvolutional2023], which extend the embedding principle to layer-wise message passing on graphs, and attention mechanisms [huangSignedGraphAttention2019, yanMUSEMultifacetedAttention2021], which adaptively weight neighbours when aggregating information. Beyond these, other neural architectures have also been explored: autoencoders [shenDeepNetworkEmbedding2020], generative adversarial networks [chakrabortySigGANAdversarialModel2022], and multilayer perceptron-based GNNs [heSSSNETSemiSupervisedSigned2022]. Other works have also targeted more complex settings, including directed signed networks [huangSDGNNLearningNode2023], bipartite structures [huangSignedBipartiteGraph2021], and temporal networks [sharmaRepresentationLearningContinuousTime2023].

### v2

@@ 4 Network structure and inference
Historically, the approach to the study of networks has been two-fold: on the one hand, we are interested in investigating their complex structures, while on the other hand, we might want to study their evolution over time and the dynamical processes occurring on them [41]. These two aspects are strongly interconnected, as the structure of a network can influence its dynamics (e.g., opinions will spread differently in a highly clustered network than in a non-clustered one) and vice versa (e.g., a change in opinion might break a link or alter its sign).
@@ 4.1 Null models
In signed networks, the most common use of null models is to detect the presence of strong or weak structural balance in empirical networks. Real-world networks often deviate from a state of perfect balance; therefore, a statistical analysis is required to determine the significance of specific patterns compared to a null model. Part of the debate over whether social systems exhibit strong or weak structural balance comes from conflicting results caused by differences in the null models employed [258, 212]. In what follows, we present null models that have mostly been developed for assessing balance. However, null models also play an important role in other contexts, for instance, to predict the creation of enmities, detect communities, or assess the importance of homophily and heterophily in signed networks, and can be used in such cases, beyond structural balance. For additional details on how to choose appropriate null models, see [179].
Figure 9: Null models for signed networks. The figure shows the different null models that can be used in signed networks, depending on the constraints that are preserved. The top row shows simple re-wiring or sign shuffle null models, while the bottom row depicts two models that preserve the signed degree. The asterisk (*) indicates that the signed degree in the STP model is preserved on average. Models in the right column preserve the underlying topology. All models preserve the ratio of negative edges in the network. This figure is an adaptation of Figure 2 from [167].
Following the classification presented in [167], null models can be organised by the constraints they preserve, such as topology, node degree, or signed degree. Figure 9 illustrates examples of the different null models presented below. In addition to the classification by Hao et al. [167], we also consider null models based on conditional probability methods, stochastic block models, and maximum-entropy formulations.
@@
This model disrupts topology while maintaining the underlying degree (i.e., the number of neighbours of each node, regardless of sign). One way to implement this null model is to randomly select two edges and interchange their source or target nodes with equal probability. To avoid multi-edges, this change is only accepted if the newly defined edge does not exist in the network. While this model is widely used in unsigned networks because it preserves node degree, it is less suitable for signed networks, because it disrupts the heterogeneity of nodes’ signed degrees and may conceal patterns related to the distribution of positive and negative links.
Sign shuffle model [370, 232]
@@
Here, the topology remains fixed while the signs of the edges are randomly redistributed. An analogous approach, shown to provide similar results in terms of balance [113], is to draw the edge signs as independent and identically distributed variables from a Bernoulli distribution. While this model preserves the total ratio of positive versus negative edges, it assumes a homogeneity in the individual ratios of each node (i.e., it assumes all nodes have the same proportion of positive edges). This assumption is dangerous because in most real-life networks there is a clear heterogeneity between individuals, some of them being more “friendly” than others [167] (see Section 5.2.3). Moreover, in [113] it is discussed how signed degree homogeneity generates more unbalanced networks and thus results in conclusions in favour of balance theory when comparing real data with such models. The explanation for this is that nodes with a lot of enmities are easier to “cast out” than uniformly distributed conflicts in the system. Another issue with this null model, brought forward by [118], is that it assumes that positive and negative edges are interchangeable, therefore testing a stronger statement than just balance. To solve this issue, the authors propose a modified permutation-based null model that stratifies the uniform permutation to edges with the same level of embeddedness.
Signed re-wire model [234]
@@
This model follows the same logic as its unsigned counterpart, explained above. However, to maintain the signed degree, it rewires the positive and negative edges separately. An equivalent implementation is to “cut” all edges in half and reattach them randomly, as long as the two halves that are reattached have the same sign [^17] In the case of a directed network, the reattachment must also follow the direction of the edge.. This approach was used, for example, in the evaluation of balance in ecological networks [338]. An alternative model, proposed in [84], is the signed Chung-Lu model, which preserves the degrees of the nodes on average rather than exactly. The authors employ this model to generate synthetic signed networks that reproduce both the global sign distribution and the proportion of balanced triangles observed in real data.
Signed Degree and Topology Preserving (STP) Model [167]
@@
Hao et al. [167] introduced a model that preserves both the network topology and the signed node degree. It is based on the maximum entropy framework [364] and provides the exact probabilities of occurrence of random graphs with both constraints. Note that, while it does preserve the signed node degree on average, it does not preserve the exact, individual signed degrees for each node.
Conditional-probability approach [230]
@@
Lerner et al. [230] proposed a different approach in terms of conditional probabilities, specifically the probability of a link having a particular sign given the presence of the link itself, instead of marginal probabilities. This approach, while it has a different implementation, has a similar conceptualisation to the topology-preserving models. The decomposition of the marginal probability of a negative link, then, would be given by:
@@
Stochastic block models [202, 404, 139]
A common null model for community detection is the stochastic block model (SBM). It assumes that nodes can be grouped into latent blocks and that the probability of observing a positive or negative edge depends only on the blocks of the endpoints. The model parameters $p^{+}_{rs}$ and $p^{-}_{rs}$ form a block matrix that specifies the probabilities of positive and negative links between nodes belonging to blocks $r$ and $s$, respectively. Jiang et al. [202] proposed a generalised signed SBM where they differentiate positive and negative block matrices. Yang et al. [404, 402] then developed a Bayesian signed SBM which explicitly accounted for edge density and frustration in the generative model. Finally, Gallo et al. [139] derived the signed SBM from an exponential random graph framework and connected it to the theory of relaxed balance (see Section 3.4).
Maximum-entropy approaches [138, 34, 166, 140]
Gallo et al. [138] have extended exponential random graph models to signed networks with both global and local constraints. Their global Signed Configuration Model with Fixed Topology (SCM-FT) imposes the topology as a hard constraint, while the positive and negative degrees of each node are matched only on average, as in the STP model explained above. The authors also define three additional variants, aligned with the types of constraints explained above: a homogeneous model with varying topology (Signed Random Graph Model or SRGM), with the same constraints as the Re-wire model, a homogeneous model with fixed topology (SRGM -FT), similar to the Sign shuffle model, and a heterogeneous model with varying topology (SCM), similar to the Signed re-wire model. A bipartite version of the SCM can also be recovered from the Bipartite Score Configuration Model in [34], and extensions for directed signed networks can be found in [166] and [140].
@@
The choice of a null model strongly shapes the interpretation of structural balance and other patterns in signed networks. In [167], the authors argue that matching the signed degrees is a relevant and often forgotten feature, while [230] shows that results may support or contradict a hypothesis of balance depending on whether conditional or marginal probabilities of edge signs are considered. Similarly, [138] finds that conclusions about strong or weak balance depend on whether heterogeneous sign propensities are included and on whether the network topology is preserved.
Beyond testing balance, null models serve broader purposes such as identifying mesoscale organisation, evaluating link or sign prediction, or constructing bipartite projections [137]. Their design should always reflect the constraints of the system under study. Re-wire models are useful in scenarios where individuals are free to decide who and how to interact with other individuals (e.g., an unconstrained social system like social media). If users’ friendliness and hostility are homogeneous, it is safe to use models that do not preserve signed degrees. However, if there are different tendencies in the way users interact with each other, the signed re-wire model would be more appropriate. Conversely, if the individuals cannot choose who they interact with but they choose the how (e.g., a high school), we would prefer topology-preserving models like a sign shuffle model, or models such as STP or SCM-FT. Ideally, one can go a few steps further and design a tailored null model when information on the allowed connections is available, since generally the topology of a network is not completely arbitrary nor strongly determined. In that case, these constraints could be incorporated in maximum-entropy-based models similar to the STP or SCM-FT [138].
@@ 4.2 Node embeddings
Node embedding methods are a powerful branch of the representation learning field, used to learn compact, low-dimensional representations of nodes, such that proximity in the embedding space reflects the underlying network structure. In this section, we review the methods specifically developed for signed networks, with a summary provided in Table LABEL:tab:node-embeddings. For a more complete overview of the field, we refer to a recent survey [416].
The core principle underlying many of these methods is synthesised in the title of one of the first papers in this domain: When Everybody Can Sit Closer to Friends than Enemies [208]. The goal of these methods is to embed the nodes in a space so that the distance between positively connected nodes (friends) is smaller than between negatively connected nodes (enemies). However, in general graphs, this ideal positioning cannot be achieved, as Kermarrec and Thraves [208] proved even for simple graphs. Consequently, most algorithms relax this objective and instead aim to minimise the number of violated constraints by optimising equivalent spectral, probabilistic, or geometric criteria. Early work in this domain focused on Euclidean and spectral approaches, while more recent contributions have leveraged advances in natural language processing, probabilistic modelling, hyperbolic geometry, and graph neural networks (GNNs).
One of the first contributions to node embeddings for signed networks was proposed by Kunegis et al. [224], who introduced a spectral approach to represent nodes in a $k$-dimensional Euclidean space. In a nutshell, the eigenvectors corresponding to the $k$ smallest eigenvalues of the opposing Laplacian are used to embed nodes in $\mathbb{R}^{k}$, so that Euclidean proximity reflects many positive links and few negative ones. This spectral embedding is particularly useful for community detection in signed graphs, as we discuss in Section 4.4. A related Euclidean embedding proposed in [92] leverages the fact that the communicability distance is Euclidean to embed the nodes within a hypersphere.
Another spectral embedding technique, inspired by physical arguments, is the SHEEP algorithm by Babul and Lambiotte [28]. The method is based on a Hamiltonian where positive edges act as springs pulling nodes together and negative edges act as repulsive forces.
@@
Minimising this Hamiltonian under a spherical constraint reduces to an eigenvalue problem involving the repelling Laplacian (Eq. 2.3). The method has been successfully applied to empirical data, including detecting opinion polarisation in online social networks [53].
A different family of methods takes inspiration from word embedding techniques like Mikolov’s word2vec model [271] and adapts them to networks. In word2vec, words are mapped to vectors by scanning sentences with sliding windows and adjusting the vectors so that words appearing in similar contexts lie close together. Grover and Leskovec adapted this idea to graphs with node2vec [156], replacing sentences by random walks on the network and words by nodes. Nodes that frequently co-occur along walks thus become close in the embedding space. The first methods to extend this principle to signed graphs were SNE [409] and Sign2vec  [198]. In these works, walks are adapted to incorporate edge signs, so that the learned vectors reflect both proximity and polarity. More recently, the stance embeddings model (SEM[314] added edge attributes, enabling embeddings that account for topical differences. Alternatively, one can map the signed network into an unsigned one without losing the information carried by negative edges. This approach, used in pipelines like ROSE [200], allows the direct use of methods such as node2vec while retaining the information contained in the negative edges in the embedding. Other approaches to signed network embedding are based on correlations between signed random walks [192], Bayesian approaches [259], hyperbolic geometry [360, 102], or latent distance models that structure embeddings within polytopes [284, 285].
The introduction of Graph Neural Networks (GNNs) opened a new phase for representation learning on signed graphs and extended node-embedding ideas into deep models. One of the earliest attempts in this direction is SiNE [393], a feedforward neural network that learns embeddings by enforcing friends to be closer than foes. This line of work has evolved toward graph convolutional neural networks [86, 210], which extend the embedding principle to layer-wise message passing on graphs, and attention mechanisms [190, 401], which adaptively weight neighbours when aggregating information. Beyond these, other neural architectures have also been explored: autoencoders [352], generative adversarial networks [60], and multilayer perceptron-based GNNs [175]. Other works have also targeted more complex settings, including directed signed networks [191], bipartite structures [189], and temporal networks [350].